### v2

@@ 3.2 Linear Complexity Bidirectional Attention
Different from the vanilla RWKV [35], we make the following modifications to its original attention mechanism to adapt it to vision tasks: (1) Bidirectional attention: We extend the upper limit of original RWKV attention from $t$ (the current token) to $T-1$ (the last token) in the summation formula to ensure that all tokens are mutually visible in the calculation of each result. Thus, the original causal attention transforms into bidirectional global attention. (2) Relative bias: We compute the absolute value of the time difference $t-i$ and divide it by the total number of tokens (denoted as $T$) to represent the relative bias of tokens in images of different sizes. (3) Flexible decay: We no longer restrict the learnable decay parameter $w$ to be positive in the exponential term allowing the exponential decay attention to focus on tokens further away from the current token in different channels. The simple yet necessary modification achieves global attention calculation and maximizes the preservation of RWKV’s low complexity and adaptability to vision tasks.
Similar to the attention in RWKV, our bidirectional attention can also be equivalently expressed in a summation form (for the sake of clarity) and an RNN form (in practical implementation).
Summation Form. The attention calculation result for the $t$-th token is given by the following formula:
@@
RNN Form. In the practical implementation, the above Eq. 5 can be transformed into a recursive formula in the form of RNN that the result of each token can be obtained through a fixed number of FLOPs. By splitting the summation term of the denominator and numerator in Eq. 5 with $t$ as the boundary, we can obtain 4 hidden states:
@@
that can be recursively computed. The update of hidden states only requires adding or subtracting one summation term and multiplying or dividing $e^{-w/T}$. Then the $t$-th result can be given as:
@@
$\text{FLOPs}(\text{Bi\mbox{-}WKV}(K,V))=13\times T\times C.\vspace{-0.5em}$ (8)
Here, the number 13 is approximately from the updates of $(a,b,c,d)$, the computation of the exponential, and the calculation of $wkv_{t}$. $T$ is the total number of update steps and is equal to the number of image tokens. The above approximation shows that the complexity of the forward process is $O(TC)$. The backward propagation of the operator can still be represented as a more complex RNN form, with a computational complexity of $O(TC)$. The specific formula for backward propagation is provided in the Appendix.
3.3 Quad-Directional Token Shift
By introducing an exponential decay mechanism, the complexity of global attention can be reduced from quadratic to linear, greatly enhancing the computational efficiency of the model in high-resolution images. However, the one-dimensional decay does not align with the neighboring relationships in two-dimensional images. Therefore, we introduce a quad-directional token shift (Q-Shift) in the first step of each spatial-mix and channel-mix module. The Q-Shift operation allows all tokens shifted and linearly interpolated with their neighboring tokens as follows:
${\rm Q\mbox{-}Shift}_{(*)}(X)$ $=X+(1-\mu_{\text{(*)}})X^{\dagger},$ (9)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ X[h,w-1,C/2:3C/4],X[h,w+1,3C/4:C]).$
Subscript (*){$(*)\in\{R,K,V\}$ denotes 3 interpolation of $X$ and $X^{\dagger}$ controlled by the learnable vectors $\mu_{(*)}$ for the later calculation of $R,K,V$, respectively. $h$ and $w$ denote the row and column index of token $X$, “:” is a slicing operation excluded the end index. The Q-Shift makes the attention mechanism of different channels obtain the prior of focusing on neighboring tokens internally without introducing many additional FLOPs. The Q-Shift operation also increases the receptive field of each token which greatly enhances the coverage of the token in the posterior layers.
3.4 Scale Up Stability
Both the number of model layers increasing and the accumulation in the exponent term during the recursion can lead to instability in the model output and affect the stability of the training process. To mitigate the instability, we employ two simple but effective modifications to stabilize the scale-up of the model. (1) Bounded exponential: As the input resolution increases, both exponential decay and growth quickly exceed the range of floating-point numbers. Therefore, we divide the exponential term by the number of tokens (such as $\mathrm{exp}(-(|t-i|-1)/T\cdot w$)), making the maximum decay and growth bounded. (2) Extra layer normalization: When the model gets deeper, we directly add layer normalization [2] after the attention mechanism and Squared ReLU operation to prevent the model’s output from overflowing. The two modifications enable stable scaling of both input resolution and model depth, allowing large models to train and converge stably. We also introduce layer scale [50] which contributes to the stability of the models as they scale up.
@@ 3.4 Scale Up Stability
Table 1: Default settings for Vision-RWKV of different scales. We report the embedding dimension, hidden dimension, and model depth for VRWKV-T/S/B/L. “Extra Norm” means additional layer normalization layers are used to stabilize the model’s outputs. “#Param” denotes the number of parameters.
3.5 Model Details
Following ViT, the hyper-parameters for variants of VRWKV, including embedding dimension, hidden dimension in linear projection, and depth, are specified in Tab. 1. Due to the increased depth of the VRWKV-L model, additional layer normalizations as discussed in Sec. 3.4, are incorporated at appropriate positions to ensure output stability.
4 Experiments
We comprehensively evaluate the substitutability of our VRWKV method for ViT in performance, scalability, flexibility, and efficiency. We validate the effectiveness of our model on the widely-used image classification dataset ImageNet [8]. For downstream dense prediction tasks, we select detection tasks on the COCO [27] dataset and semantic segmentation on the ADE20K [62] dataset.
@@ 4 Experiments
hierarchical ResNet-18 [20] $224^{2}$ 11.7M 1.8G 69.9
PVT-T [56] $224^{2}$ 13.2M 1.9G 75.1
ResNet-50 [20] $224^{2}$ 25.6M 4.1G 76.6
Swin-T [30] $224^{2}$ 28.3M 4.4G 81.2
PVT-M [56] $224^{2}$ 44.2M 6.7G 81.2
ResNet-101 [20] $224^{2}$ 44.6M 7.9G 78.0
Swin-S [30] $224^{2}$ 49.6M 8.7G 83.0
PVT-L [56] $224^{2}$ 61.4M 9.8G 81.7
Swin-B [30] $224^{2}$ 87.8M 15.1G 83.4
DeiT-T [49] $224^{2}$ 5.7M 1.3G 72.2
DeiT-S [49] $224^{2}$ 22.1M 4.6G 79.9
XCiT-S12 [1] $224^{2}$ 26.0M 4.8G 82.0
DeiT-B [49] $224^{2}$ 86.6M 17.6G 81.8
XCiT-L24 [1] $224^{2}$ 189.0M 36.1G 82.9
ViT-L [12] $384^{2}$ 309.5M 191.1G 85.2
@@
non-hierarchical ${}^{\star}$ $384^{2}$ 334.9M 189.5G 86.5
Table 2: Validation results on ImageNet-1K. VRWKV-T/S/B are trained from scratch using ImageNet-1K, while VRWKV-L is pre-trained on Imagenet-22K and fine-tuned on Imagenet-1K. “#Param” denotes the number of parameters, and “FLOPs” means the computational workload for processing an image at the resolution specified in the “Size” column. “normal-$\star$” denotes Bamboo-47K [61] is used in the pre-training.
@@ 4.1 Image Classification
Settings. For -Tiny/Small/Base models, we conduct supervised training from scratch on ImageNet-1K [8]. Following the training strategy and data augmentation of DeiT [49], we use a batch size of 1024, AdamW [33] with a base learning rate of 5e-4, weight decay of 0.05, and cosine annealing schedule [32]. Images are cropped to the resolution of $224\times 224$ for training and validation. For the -Large models, we first pre-train them for 90 epochs on ImageNet-22K with a batch size of 4096 and resolution of $192\times 192$, and then fine-tune them for 20 epochs on ImageNet-1K to a higher resolution of $384\times 384$.
Results. We compare the results of our VRWKV with other hierarchical and non-hierarchical backbones on the ImageNet-1K validation dataset. As shown in Tab. 2, with the same number of parameters, computational complexity, and training/testing resolutions, VRWKV achieves better results than ViT. For example, when VRWKV-T has slightly lower FLOPs than DeiT-T(1.2G vs. 1.3G), our VRWKV-T achieves a 2.9 point higher than DeiT-T on top-1 accuracy. When the model size scales up, VRWKV still demonstrates higher baseline performance. In the case of large models, VRWKV-L achieves a 0.8 point higher top-1 accuracy of 86.0% at the resolution of $384\times 384$ than ViT-L, with a slightly reduced computational cost. The superior performance from tiny to large-size models demonstrates that the VRWKV model possesses the scalability as ViT. Additionally, after using Bamboo-47K [61] in the pre-train process, the performance of VRWKV-L can be further boosted to 86.5%, indicating that our VRWKV also possesses the ability like ViT to benefit from pre-training on large-scale datasets. The exploration of VRWKV in classification tasks demonstrates its potential to be a viable alternative to traditional ViT models.
@@ 4.2 Object Detection
Method #Param FLOPs $\rm AP^{b}$ $\rm AP^{m}$
${}^{\dagger}$ [49] 8.0M 95.4G 41.1 37.5
ViT-T [49] 8.0M 147.1G 41.6 37.9
@@
Table 3: Object detection and instance segmentation on COCO val2017. All models adopt the ViT-Adapter [4] to generate multi-scale features for detection heads.  -T/S/B models are initialized with ImageNet-1K pre-trained weights, and all -L models with ImageNet-22K weights. “#Param” denotes the number of backbone parameters. “FLOPs” denotes the computational workload of the backbone with an input image of $1333\times 800$. “normal-$\dagger$” means window attention is adopted in ViT layers.
Settings. In the detection tasks, we adopt Mask R-CNN [19] as the detection head. For the -Tiny/Small/Base models, the backbones use weights pre-trained on ImageNet-1K for 300 epochs. For the -Large model, weights pre-trained on ImageNet-22K are used. All models use a 1$\times$ training schedule (i.e., 12 epochs) with a batch size of 16, and AdamW [33] optimizer with an initial learning rate of 1e-4 and weight decay of 0.05.
Results. In Tab. 3, we report the detection results on the COCO val [27] dataset using VRWKV and ViT as backbones. As the results showed in Fig. 1 (a) and Tab. 3, due to the use of window attention in dense prediction tasks, VRWKV with global attention can achieve better performance than ViT with lower FLOPs. For example, VRWKV-T has approximately 30% lower backbone FLOPs compared to ${}^{\dagger}$, with an improvement of $\mathrm{AP^{b}}$ by 0.6 points. Similarly, VRWKV-L achieves a 1.9-point increase in $\mathrm{AP^{b}}$ with lower FLOPs compared to ${}^{\dagger}$. Additionally, we compare the performance of VRWKV and ViT using global attention. For instance, VRWKV-S achieves similar performance to ViT-S with 45% lower FLOPs. This demonstrates the effectiveness of VRWKV’s global attention mechanism in dense prediction tasks and the advantage of lower computational complexity compared to the original attention mechanism.
@@ 4.3 Semantic Segmentation
Method #Param FLOPs mIoU
ViT-T [49] 8.0M 20.9G 42.6
VRWKV-T (ours) 8.4M 16.6G 43.3
ViT-S [49] 27.5M 54.0G 46.2
VRWKV-S (ours) 29.3M 46.3G 47.2
ViT-B [49] 99.5M 157.9G 48.8
VRWKV-B (ours) 106.6M 146.0G 49.2
ViT-L [44] 327.0M 446.8G 53.4
VRWKV-L (ours) 351.9M 421.9G 53.5
Table 4: Semantic segmentation on the ADE20K val set. All models used ViT-Adapter [4] for multi-scale feature generation and are trained with UperNet [59] as the segmentation heads. For consistency in comparison, all -T/S/B models are initialized using ImageNet-1K pre-training, whereas -L models utilize ImageNet-22K pre-training. “#Param” refers to the number of parameters of the backbone. We report the FLOPs of backbones with the input size of $512\times 512$.
Settings. In the semantic segmentation task, we use UperNet [59] as the segmentation head. Specifically, all ViT models use global attention in the segmentation task. For the -Tiny/Small/Base models, the backbones use weights pre-trained on ImageNet-1K. And for the -Large model, weights pre-trained on ImageNet-22K are used. We employ the AdamW optimizer with an initial learning rate of 6e-5 for the -Small/Base/Large models and 12e-5 for the -Tiny model, a batch size of 16, and a weight decay of 0.01. All models are trained for 160k iterations on the training set of the ADE20K dataset [62].
Results. As shown in Tab. 4, when using UperNet for semantic segmentation, models based on VRWKV consistently outperform those based on ViT with global attention, while also being more efficient. For example, VRWKV-S achieves 1 point higher than ViT-S with a 14% FLOPs decrease. VRWKV-L creates a result of 53.5 mIoU similar to ViT-L while the computation of the backbone has a 25G FLOPs lower. These results demonstrate that our VRWKV backbones can extract better features for semantic segmentation compared to ViT backbones while also being more efficient, benefiting from the linear complexity attention mechanism.

### v3

@@ 3.2 Linear Complexity Bidirectional Attention
Different from the vanilla RWKV (Peng et al., 2023), we make the following modifications to its original attention mechanism: (1) Bidirectional attention: We extend the upper limit of original RWKV attention from $t$ (the current token) to $T-1$ (the last token) to ensure that all tokens are mutually visible in the calculation of each result. Thus, the original causal attention transforms into bidirectional global attention. (2) Relative bias: We compute the absolute value of the time difference $t-i$ and divide it by the total number of tokens (denoted as $T$) to represent the relative bias of tokens in images of different sizes. (3) Flexible decay: We no longer restrict the learnable decay parameter $w$ to be positive in the exponential term allowing the exponential decay attention to focus on tokens further away from the current token.
Under the collective influence of these ingenious modifications, we achieve global attention while maintaining linear complexity to the input token number $T$, thereby maximizing the preservation of RWKV’s inherent low complexity and extending it to the visual domain.
Similar to RWKV, our bidirectional attention can also be equivalently expressed in a summation form (for the sake of clarity) and an RNN form (in practical implementation).
Summation Form. The attention calculation result for the $t$-th token is given by the formula:
@@
RNN Form. In the practical implementation, the above Eq 5 can be transformed into a recursive formula in the form of RNN that the result of each token can be obtained through a fixed number of FLOPs. By splitting the summation term of the denominator and numerator in Eq 5 with $t$ as the boundary, we can obtain 4 hidden states:
@@
that can be recursively computed due to its mathematical formulation. The update of hidden states only requires adding or subtracting one summation term and multiplying or dividing $e^{-w/T}$. Then the $t$-th result can be given as:
@@
$\text{FLOPs}(\text{Bi\mbox{-}WKV}(K,V))=13\times T\times C.\vspace{-1em}$ (8)
Here, the number 13 is approximately from the updates of $(a,b,c,d)$, the computation of the exponential, and the calculation of $wkv_{t}$. The above approximation shows that the complexity of the forward process is $O(TC)$. The backward propagation of the operator can still be represented as a more complex RNN form, with a computational complexity of $O(TC)$. The specific formula for forward updating and backward propagation is provided in the Appendix A.1.
@@
Table 1: Default settings for Vision-RWKV of different scales. We report the embedding dimension, hidden dimension, and model depth. “Extra Norm” means additional layer normalization layers are used to stabilize the model’s outputs. “#Param” denotes the number of parameters.
3.3 Quad-Directional Token Shift
We introduce a quad-directional token shift (Q-Shift) as a flexible extension of the original token shift operation in RWKV in the first step of each spatial-mix and channel-mix module. The Q-Shift operation allows all tokens shifted and linearly interpolated with their neighboring tokens as follows:
${\rm Q\mbox{-}Shift}_{(*)}(X)$ $=X+(1-\mu_{\text{(*)}})X^{\dagger},$ (9)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ X[h,w-1,C/2:3C/4],X[h,w+1,3C/4:C]).$
Subscript (){$(*)\in\{R,K,V\}$ denotes 3 interpolation of $X$ and $X^{\dagger}$ controlled by the learnable vectors $\mu_{(*)}$ for the later calculation of $R,K,V$, respectively. $h$ and $w$ denote the row and column index of token $X$, “:” is a slicing operation excluded the end index. The Q-Shift makes the attention mechanism of different channels obtain the prior of focusing on neighboring tokens internally without introducing many additional FLOPs. It also increases the receptive field of each token which greatly enhances the coverage of the token in the posterior layers. $X^{\dagger}$ is obtained by slicing X without introducing new computations, allowing for flexible transformations during the training process for different tasks. When handling sparse inputs that do not contain original image space information, such as masked image modeling, shifting can be applied in a single dimension to maximize the preservation of image priors.
3.4 Scale Up Stability
Increasing model depth and the accumulation of exponential terms during recursion can lead to instability in the training process. To mitigate this, we propose two simple yet effective adjustments: (1) Bounded exponential: As input resolution increases, both exponential decay and growth can quickly exceed the range of floating-point numbers. To address this, we divide the exponential term by the number of tokens (e.g., $\mathrm{exp}(-(|t-i|-1)/T\cdot w$)), making the maximum decay and growth bounded. (2) Extra layer normalization: As models become deeper, we apply layer normalization (Ba et al., 2016) after the attention mechanism and the Squared ReLU operation, to prevent the model’s output from overflowing. These two adjustments promote stable scaling of input resolution and model depth, facilitating the smooth convergence of large models. Additionally, we incorporate layer scale (Touvron et al., 2021b), which further enhances model stability during scaling.
@@ 3.4 Scale Up Stability
hierarchical ResNet-18 (He et al., 2016) $224^{2}$ 11.7M 1.8G 69.9
PVT-T (Wang et al., 2021) $224^{2}$ 13.2M 1.9G 75.1
ResNet-50 (He et al., 2016) $224^{2}$ 25.6M 4.1G 76.6
Swin-T (Liu et al., 2021) $224^{2}$ 28.3M 4.4G 81.2
PVT-M (Wang et al., 2021) $224^{2}$ 44.2M 6.7G 81.2
ResNet-101 (He et al., 2016) $224^{2}$ 44.6M 7.9G 78.0
Swin-S (Liu et al., 2021) $224^{2}$ 49.6M 8.7G 83.0
PVT-L (Wang et al., 2021) $224^{2}$ 61.4M 9.8G 81.7
Swin-B (Liu et al., 2021) $224^{2}$ 87.8M 15.1G 83.4
DeiT-T (Touvron et al., 2021a) $224^{2}$ 5.7M 1.3G 72.2
DeiT-S (Touvron et al., 2021a) $224^{2}$ 22.1M 4.6G 79.9
XCiT-S12 (Ali et al., 2021) $224^{2}$ 26.0M 4.8G 82.0
DeiT-B (Touvron et al., 2021a) $224^{2}$ 86.6M 17.6G 81.8
XCiT-L24 (Ali et al., 2021) $224^{2}$ 189.0M 36.1G 82.9
ViT-L (Dosovitskiy et al., 2020) $384^{2}$ 309.5M 191.1G 85.2
@@
VRWKV-L $384^{2}$ 334.9M 189.5G 86.2
non-hierarchical VRWKV-L $384^{2}$ 334.9M 189.5G 86.5
Table 2: Validation results on ImageNet-1K. VRWKV-T/S/B are trained on ImageNet-1K, while VRWKV-L is pre-trained on Imagenet-22K and fine-tuned on Imagenet-1K. “#Param” denotes the number of parameters, and “FLOPs” represents the computational workload for the specified image resolution in the “Size” column. $\dagger$” means additional MAE pre-training is applied in the pre-training process. $\star$” indicates Bamboo-47K is used in the pre-training.
3.5 Model Details
Following ViT, the hyper-parameters for variants of VRWKV, including embedding dimension, hidden dimension in linear projection, and depth, are specified in Table 1. Due to the increased depth of the VRWKV-L model, additional layer normalizations as discussed in Section 3.4, are incorporated at appropriate positions to ensure output stability.
4 Experiments
We comprehensively evaluate the substitutability of our VRWKV for ViT in performance, scalability, flexibility, and efficiency. The model’s effectiveness is validated in image classification, object detection, and semantic segmentation tasks.
@@ 4.1 Image Classification
Settings. For -Tiny/Small/Base models, we conduct supervised training from scratch on ImageNet-1K (Deng et al., 2009). Following the training strategy and data augmentation of DeiT (Touvron et al., 2021a), we use a batch size of 1024, AdamW (Loshchilov & Hutter, 2017) with a base learning rate of 5e-4, weight decay of 0.05, and cosine annealing schedule (Loshchilov & Hutter, 2016). Images are cropped to the resolution of $224\times 224$ for training and validation. For the -Large models, we first pre-train them for 90 epochs on ImageNet-22K with a batch size of 4096 and resolution of $192\times 192$, and then fine-tune them for 20 epochs on ImageNet-1K to a higher resolution of $384\times 384$.
Results. We compare the results of our VRWKV with other hierarchical and non-hierarchical backbones on the ImageNet-1K validation dataset. As shown in Table 2, with the same number of parameters, computational complexity, and training/testing resolutions, VRWKV achieves better results than ViT. For example, when VRWKV-T has slightly lower FLOPs than DeiT-T(1.2G vs 1.3G), our VRWKV-T achieves a 2.9 point higher than DeiT-T on top-1 accuracy. When the model size scales up, VRWKV still demonstrates higher baseline performance. VRWKV-L achieves a 0.8 point higher top-1 accuracy of 86.0% at the resolution of $384\times 384$ than ViT-L, with a slightly reduced computational cost. The superior performance from tiny to large-size models demonstrates that the VRWKV model possesses the scalability as ViT. Additionally, after using a larger dataset Bamboo-47K (Zhang et al., 2022) in the pre-train process, the performance of VRWKV-L can be further boosted to 86.5%, indicating that our VRWKV also possesses the ability like ViT to benefit from pre-training on large-scale datasets. The exploration of VRWKV in classification tasks demonstrates its potential to be a viable alternative to traditional ViT models.
@@ 4.2 Object Detection
Method Window #Param FLOPs $\rm AP^{b}$ $\rm AP^{m}$
ViT-T (Touvron et al., 2021a) 8.0M 95.4G 41.1 37.5
ViT-T (Touvron et al., 2021a) 8.0M 147.1G 41.6 37.9
@@
Table 3: Object detection and instance segmentation on COCO val2017. All models adopt the ViT-Adapter (Chen et al., 2023) to generate multi-scale features for detection heads.  -T/S/B models are initialized with ImageNet-1K weights, and all -L models use ImageNet-22K weights. “#Param” indicates the backbone parameter, and ’FLOPs’ represent the backbone’s computational workload for a $1333\times 800$ input. “Window” denotes the use of window operation in ViT layers.
Settings. In the detection tasks, we adopt Mask R-CNN (He et al., 2017) as the detection head. For the -Tiny/Small/Base models, the backbones use weights pre-trained on ImageNet-1K for 300 epochs. For the -Large model, weights pre-trained on ImageNet-22K are used. All models use a 1$\times$ training schedule (i.e., 12 epochs) with a batch size of 16, and AdamW (Loshchilov & Hutter, 2017) optimizer with an initial learning rate of 1e-4 and weight decay of 0.05.
Results. In Table 3, we report the detection results on the COCO val (Lin et al., 2014) dataset using VRWKV and ViT as backbones. As the results showed in Figure 1(a) and Table 3, due to the use of window attention in dense prediction tasks, VRWKV with global attention can achieve better performance than ViT with lower FLOPs. For example, VRWKV-T has approximately 30% lower backbone FLOPs compared to ViT-T using window attention, with an improvement of $\mathrm{AP^{b}}$ by 0.6 points. Additionally, we compare the performance of VRWKV and ViT using global attention. For instance, VRWKV-S achieves similar performance to ViT-S with 45% lower FLOPs. This demonstrates the effectiveness of VRWKV’s global attention mechanism in dense prediction tasks and the advantage of lower computational complexity compared to the original attention mechanism.
Method Window #Param FLOPs mIoU
ViT-T (Touvron et al., 2021a) 8.0M 20.9G 42.6
VRWKV-T (ours) 8.4M 16.6G 43.3
ViT-S (Touvron et al., 2021a) 27.5M 54.0G 46.2
VRWKV-S (ours) 29.3M 46.3G 47.2
ViT-B (Touvron et al., 2021a) 99.5M 157.9G 48.8
VRWKV-B (ours) 106.6M 146.0G 49.2
ViT-L (Steiner et al., 2021) 327.0M 446.8G 53.4
VRWKV-L (ours) 351.9M 421.9G 53.5
Table 4: Semantic segmentation on the ADE20K val set. All models used ViT-Adapter (Chen et al., 2023) for multi-scale feature generation and are trained with UperNet as the segmentation heads. For consistency in comparison, all -T/S/B models are initialized using ImageNet-1K pre-training, whereas -L models utilize ImageNet-22K pre-training. “#Param” refers to the number of parameters of the backbone. We report the FLOPs of backbones with the input size of $512\times 512$.
@@ 4.3 Semantic Segmentation
Settings. In the semantic segmentation task, we use UperNet (Xiao et al., 2018) as the segmentation head. Specifically, all ViT models use global attention in the segmentation task. For the -Tiny/Small/Base models, the backbones use weights pre-trained on ImageNet-1K. And for the -Large model, weights pre-trained on ImageNet-22K are used. We employ the AdamW optimizer with an initial learning rate of 6e-5 for the -Small/Base/Large models and 12e-5 for the -Tiny model, a batch size of 16, and a weight decay of 0.01. All models are trained for 160k iterations on the training set of the ADE20K dataset (Zhou et al., 2017).
Results. As shown in Table 4, when using UperNet for semantic segmentation, models based on VRWKV consistently outperform those based on ViT with global attention, while also being more efficient. For example, VRWKV-S achieves 1 point higher than ViT-S with a 14% FLOPs decrease. VRWKV-L creates a result of 53.5 mIoU, similar to ViT-L, while the computation of the backbone has a 25G FLOPs lower. These results demonstrate that our VRWKV backbones can extract better features for semantic segmentation compared to ViT backbones while also being more efficient, benefiting from the linear complexity attention mechanism.